OpenStudy (anonymous):

M is supplemental to an angle whose measure is 125 degrees and P is the complement of an angle whose measure is 70 degrees. What is M − P? 35 degrees 30 degrees 70 degrees 195 degrees

OpenStudy (tad1):

180 - 125 = 55 = M 90-70 = 20 = P M - P = 55 - 20= 35

OpenStudy (anonymous):

ok, thank you !(:

OpenStudy (tad1):

you are welcome!